A Markov chain {Xn, n ≥ 0} with state space S = {0, 1, 2} has transition probability matrix

0.1 0.3 0.6

p = 0.5 0.2 0.3

0.4 0.2 0.4

If P(X0 = 0) = P(X0 = 1) = 0.4 and P(X0 = 2) = 0.2, find the distribution of X2 and evaluate P[X2 < X4].
Answer #1

P(X​​​​​​​​​4 = 1/ X​​​​​​​​​2= 0) = P(X​​​​​​​​​​​2 = 1/X​​​​​​0 = 0) = (P​​​​​​2)0,1  = 0.21SOLUTIO o3ore-slep tunsihon Probability mai p(.zo) : p( % ● ! ): 。.4 , p(X.L) =o.2 o 0.3 6 6.5 0.2 0.3 0.27 025 48 0.3 28 +p(C021) (0 3 2 8) + 6. 39) Co 328 +4S) (0.232) 二 o, 3081 Ars​​,  which is element corresponding to 1 st row 2 nd column.